OpenStudy (anonymous):

You have 1.50mL of a urine specimen. The assay you must do requires a 1/25 dilution. How much water must be added to the specimen to achieve this dilution?

OpenStudy (amistre64):

urine? eww!!!

OpenStudy (amistre64):

1/25 = .04 as a decimaled percent. 1.5(1) + y(0) = (1.5+y)(.04)

OpenStudy (amistre64):

1.5 = .06 + .04y 1.5 - .06 = .04y y = 1.44/.04 = 144/4 = 36 36mL of water...and some soap?

OpenStudy (anonymous):

hahaha. thank you :)

OpenStudy (anonymous):

i did get 36 as well!

OpenStudy (amistre64):

yay!! we shuld join mensa :)